OpenStudy (anonymous):

∫∫e^(x/y) dy dx limits for 1<=y<=x & 1<=x<=0. evaluate reversing the order of integration.

OpenStudy (anonymous):

\[\int\limits_{1}^{0}\int\limits_{1}^{x}e^{\frac{x}{y}}dydx\]like this?

OpenStudy (anonymous):

the limits are in the opposite dirn

OpenStudy (anonymous):

for both integrals?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. your region of integration looks like this:We want it in the opposite order which looks like this:This gives:\[\int\limits_{0}^{1}\int\limits_{y}^{1}e^\frac{x}{y}dxdy\]